OpenStudy (anonymous):

The vectors <8,4> and <8,k> are perpendicular. Find k .

OpenStudy (amistre64):

do a dot product to solve for k

OpenStudy (amistre64):

or, swap the first one, and negate an element and scale it ...

OpenStudy (amistre64):

type in the reply box down here to respond in a post ....

OpenStudy (amistre64):

how what? dot product? or a perp property?

OpenStudy (anonymous):

any of them that's easier? what is the answer ?

OpenStudy (amistre64):

we dont give out direct answers; we try to help you approach a solution. You must have come across something in your material that relates to this. Im just assuming you have heard of a dot product; or have been exposed to the perpendicular slope property

OpenStudy (anonymous):

none, im doing the aleks prgram as a pre req for my physiqs ,

OpenStudy (anonymous):

i assume the answer would be -4 isn't it?

OpenStudy (amistre64):

i like -4 as an answer. just not too sure how you approached it :)

OpenStudy (anonymous):

according to here... http://mathworld.wolfram.com/PerpendicularVector.html

OpenStudy (amistre64):

for a dot product, I stack the vectors and multiply down the columns <8,4> <8,k> ------ 64+4k <-- then add the row perp vectors have a dot product that is equal to 0 64+4k = 0 4k = -64 k = -8 ; hmm, im liking -8 better than -4 at the moment

OpenStudy (amistre64):

pfft, i spose the test i just got out of took a toll on me ....

OpenStudy (amistre64):

-64/4 = -16 :) the perpendicular slope property would have given us this \[\frac48\frac k8=-1\] \[\frac4{64}k=-1\] \[k=-\frac{64}{4}\] as well

OpenStudy (anonymous):

ok thanks

OpenStudy (amistre64):

... now i have to worry about that test lol